# Utilizing and extending superconducting circuit toolbox to simulate analog quantum gravity

### v1

@@ Abstract
There has been considerable effort to simulate quantum gravity features in solid state systems, such as analog black or wormholes in cold atoms or tilted Weyl semimetals, or holographic aspects on quantum processors. Surprisingly however, superconducting circuits have so far received only limited attention in this regard. Moreover, for quantum superpositions of spacetime geometries – a highly contentious notion within the quantum gravity community itself – there currently exist no solid state blueprints. We here show how quantum circuit hardware can implement a variety of classical and quantum spacetime geometries on lattices, by both using established circuit elements and introducing new ones. While gyrators tilt the dispersion relation, an event horizon requires in addition negative inductances, which we propose to realize with transient flux drive of regular Josephson junctions. Already for classical wormhole simulations, our approach contains several surprising features. We demonstrate the possibility of a metric sharply changing within a single lattice point, thus entering a regime where the spacetime curvature itself is trans-Planckian, and the Hawking temperature genuinely ill-defined. In fact, our approach suggests that stable, thermal event horizons are incompatible with strictly discrete lattice models. We thus propose to directly probe the evaporation of a wormhole by tracking the accumulation of charge and phase quantum fluctuations over short time scales, which are a robust signature even in the presence of a dissipative environment. Moreover, we present a loop-hole for the typical black/while hole ambiguity in lattice simulations: the existence of exceptional points in the dispersion relation allows for the creation of pure black (or white) hole horizons – at the expense of radically changing the interior wormhole dynamics. Finally, based on multistable Josephson junctions, we introduce the notion of quantum inductors: circuit elements that can be prepared in a superposition of different inductance values. Such inductors realize regions with a quantum superposition of analog spacetime. The entanglement of signals with the quantum spacetime can be probed by a type of delayed-choice experiment. In the outlook, we provide a pathway towards more complex simulations including a quantum coherent dynamics of spacetime itself, as well as prospects and ideas for experimental verification of the basic concepts.
@@ I Introduction
Unifying quantum mechanics and gravity poses several theoretical [1, 2, 3] and experimental challenges [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], which motivates studying analogs of relativistic quantum effects in experimentally accessible systems [15]. Several ideas have therefore been proposed for simulating gravitational effects in solid-state systems—for example, holographic ideas borrowed from string theory  [16, 17, 18, 19, 20, 21], or more directly, using Unruh’s proposal of "sonic black holes" [22] as an inspiration to simulate black holes in labs. Along these lines, apparent event horizons and the resulting Hawking radiation have been studied on various platforms  [23, 24, 25, 26, 27, 28, 29, 30, 31], most notably in ultracold atomic gases [32, 33, 34, 35].
@@
There remain however a number of open questions. First, so far for all solid state implementations the metric can change only over a finite “healing length”. If the curvature at the horizon is too weak, the resulting small Hawking temperature may thus be below the threshold of the system’s intrinsic temperature [36] – an obstacle that (at least up to now) seems to have only been overcome for cold atom simulators [35]. Relatedly, systems with an apparent event horizon generically do not have a well-defined ground state, such that any coupling to the environment leads to an instability, making it extremely challenging to distinguish intrinsic radiation due to the horizon (which would actually simulate aspects of a universe with nontrivial spacetime metric) and radiation due to environment induced relaxation. In addition, even in the absence of environment, the closed system can in some platforms be spontaneously unstable [32, 33]. Furthermore, lattice realizations (which ultimately concerns all solid state systems) add a number of interesting but challenging facets, such as a natural (though possibly artifical) resolution of the trans-Planckian problem [37, 38, 39], and most notably, the fact any event horizon has both black and white hole character, since the dispersion relation in a periodic Brillouin zone must necessarily cross zero at least twice [25]. Finally, within the quantum gravity community, there also exists the highly contentious idea of superpositions of spacetime geometries, originating from attempts at quantising the metric tensor [40, 41]. In particular, recent theoretical proposals suggest experiments that examine quantum gravitational effects caused by a superposition of different mass configurations [42, 43, 44, 45, 46, 47, 48, 49]. To the best of our knowledge, there currently exist no feasible blueprints for solid state systems capable of simulating quantum superpositions of spacetime geometries.
@@
Superconducting circuits are one of the prime candidates for building large-scale quantum hardware [50, 51, 52]. The behavior of these circuits and their interactions with quantized electromagnetic fields in the microwave range are described by circuit quantum electrodynamics (cQED), a formalism that reduces circuits to lumped elements whose phase and charge are canonically conjugate [53, 54, 55, 56, 57, 58, 59, 60, 61]. The toolbox of superconducting circuits includes nonlinear elements such as Josephson junctions and linear elements such as capacitors, inductors, and gyrators—a nonreciprocal multi-port element [62, 63, 64, 65, 66, 67]. The same nonreciprocal behaviour also occurs due to topological transitions in the transport degrees of freedom of multiterminal junctions [68, 69, 70, 71, 72, 73] Over the course of decades, it has been shown that networks of superconducting circuits can give rise to a large number of physical phenomena and imitate various quantum field theories from other domains [74, 75, 76, 77, 78, 79, 80, 81, 82, 28, 83]. However, the exploration of analog gravity phenomena has so far been limited to Hawking-like radiation in solitons of Josephson junction arrays [26, 27, 84, 85, 86, 87, 28, 29] which leave only little control over the shape of the spacetime geometry, or recreation of a limited class of curved spacetimes by means of flux control of SQUID arrays [30, 31].
@@
Moreover, we find an instructive loop-hole to the aforementioned ambiguity of black- versus white-hole event horizons in lattice systems. Namely, the inductive coupling can go either via nearest or next-to-nearest neighbour nodes. For the former, the dispersion relation inside the wormhole region exhibits an exceptional point, such that the eigenspectrum no longer crosses zero energy twice, but instead takes a detour on the complex plane to satisfy periodicity. We thus realize lattice versions of event horizons which are definitely either black- or white hole, but not both. However, this feature radically changes the dynamics in the interior of the wormhole: instead of an evaporation from the horizons outwards the entire wormhole interior evaporates everywhere immediately after the quench.
@@
Finally, we go beyond the simulation of classical apparent horizons, and demonstrate that the cQED platform suggested here allows for the implementation of superposition of spacetime geometries. This idea draws from the recent development of multistable Josephson junctions [92, 93, 94], i.e., junctions whose energy-phase relationship has more than one local minimum. Specifically, the engineering of multiple minimas with different local curvature creates a type of quantum inductor – an element that can be prepared in a superposition of different inductance values. When creating a chain of maximally entangled quantum inductors, we propose to create a superposition of different spacetime geometries. A classical signal traversing this region propagates in a superposition of different velocities with respect to the lab frame, and thus entangles with the quantum spacetime. The entanglement can be probed by a delayed-choice type of experiment, where a Hadamard gate is applied on the quantum inductor array after traversal of the quantum spacetime region, leading to a specific interference pattern.
@@
In an extended outlook, we show that in particular the idea of quantum superpositions of the spacetime geometries harbors a great potential for follow-up studies, where one could endow the spacetime itself with a nonzero Hamiltonian, thus working towards much more complex simulations of analog quantum gravity, where both the field moving in curved analog spacetime and the spacetime itself have a quantum coherent and strongly correlated dynamics. Moreover, while we mostly consider long chains in the main part of this work, as they provide a clean interpretation of the effects in the context of curved spacetime, we are aware that they may be challenging to realize experimentally. In the outlook, we therefore also propose experiments on only few lattice points as proofs of principle, which already contain much of the pertinent phenomenology present in long chains.
@@
This paper is organized as follows, in section III we introduce and study the two circuits that will host the appparent horizons that we wish to investigate, we especially focus on the different effect that negative inductances have on the dispersion relations of these two circuits anticipating the differences between the horizons in them. Section IV is a brief detour, where we expand upon the idea of using flux quench on a Josephson junction to get a negative inductance, and in section V we study and characterize the horizons in the two aforementioned circuits, by looking at the time evolution of quantum fluctuations of the conjugate charge and the phase difference. Finally, in section VII we present our proposal for observing quantum superposition of two spacetime geometries using multi-stable Josephson junctions.
@@ II.2 Simulation with circuit networks
As we see in Eq. (2), for a system to emulate a field moving on a nontrivial spacetime, we need interactions that provide the terms $\sim\dot{\phi}^{2}$, $\sim\dot{\phi}\partial_{x}\phi$, and $\sim(\partial_{x}\phi)^{2}$. Quantum circuit theory [54, 56, 63, 61] readily provides elements that (in a certain continuum limit) provide such interactions. For a superconducting node $j$, a finite capacitance ($C_{j}$) provides to the Lagrangian $\mathcal{L}$ an energy contribution of the form
@@
    we show how to realize negative inductances via transient flux drive, and thus create overtilted dispersion relations and analog (apparent) event horizons.
Importantly, the setup we propose allows for a radical change of the spacetime geometry over only a few lattice sites (and as we show in a moment, even over a single lattice bond). With the energy scales of the local charging energy given by $E_{C}$ and when using Josephson junctions as inductive elements $E_{J}$, we would thus (at least in principle) be able to reach Hawking temperatures (as defined above) on the order of $\sim\sqrt{E_{C}E_{J}}$ which (taking as a typical transmon qubit frequency the value of $1\text{GHz}\sim 10\text{GHz}$) can yield up to $0.1\text{K}\sim 1\text{K}$ (which is above typical cryogenic temperatures).
@@
At any rate, some of the lattice-specific stability issues can be appreciated on the general level here, others will be shown subsequently for a specific model. On a general footing, we see in Eq. (2) that any (constant or spatially dependent) local tilt $v$ (which in our simulation would be implemented by constant or spatially varying gyrator parameters $G_{j}$) must equally appear in the $(\partial_{x}\phi)^{2}$ term, as the corresponding matrix element $g^{11}$ ($\mu,\nu=1$ is the space component) is proportional to $v^{2}-u^{2}$. In other words, a finite nonreciprocity in the circuit needs to be “countered” by a corresponding inductive interaction, in order to realize actions of the form of Eq. (2). Consequently, we would need an according spatially dependent tuning of the local inductance profile $L_{j}$ which matches exactly with the values of $G_{j}$. But first of all, such a perfect fine tuning is realistically not possible, such that the condition $g=-1$ is in general only approximately, but never exactly, fulfilled. Hence, the mapping onto Eq. (5) is likewise not exact, and we cannot guarantee a perfect causal decoupling across the horizon, and the closed system might become spontaneously unstable over already short time scales. In fact, below we refrain from a simulation with constant metric determinant $g$, such that the systems we consider do not map to Eq. (5) in the continuum limit.
@@
Thus, we see that in order to have the correct counter term in the $g^{11}$ component for the lattice, we actually do not need nearest neighbour inductive coupling (as indicated in Eq. (8)) but instead next-to-nearest neighbour coupling, $\sim(\phi_{j+1}-\phi_{j-1})^{2}$. As further detailed below, because of this fundamental feature of the gyrator element, nearest and next-to-nearest neighbour inductive couplings have fundamentally different stability properties. In particular,
@@
The breaking of charge conservation (which is physical as it corresponds to leakage to ground) effectively provides a mass term in the theory. In fact, we will include a (finite but very small) mass term in the following, but for a different reason: it allows us deal with the otherwise problematic zero mode in the boson Hamiltonian.
@@ III Setup
We here show with two example circuit models the capability of circuit arrays to engineer nontrivial dispersion relations with analog wormholes. The first proposed circuit consists of an array of $J$ nodes (with periodic boundary conditions, i.e., $J+1=1$), interconnected via Josephson junctions in parallel with 3-port gyrators, as depicted in Fig. 1a). The complete Lagrangian of the circuit will consist of three parts, namely a capacitive Lagrangian in Eq. (7), a gyrator Lagrangian in Eq. (9) and a term for the Josephson junctions $\mathcal{L}_{J}=\sum_{j=1}^{J}E_{J}\cos\left(\phi_{j+1}-\phi_{j}+\phi_{\text{ext},j}\right)$. Here, $E_{J}$ is the Josephson energy and $\phi_{\text{ext},j}$ is the external flux coupling to each Josephson junctions. This flux, and its time-dependent control, will play a pivotal role in what follows.
@@
As stated in the previous section, the stability of circuits where Josephson junctions connect nearest neighbors nodes and the ones where Josephson junctions couple next nearest neighbor nodes is very different. Therefore to illustrate their differences the second circuit we study has next nearest neighbor connections via Josephson junctions, see Fig. 1b). The Hamiltonian for such a circuit is
@@
For small $k$, the system exhibits a linear dispersion relation, see [see Figs. 2a) and 2b)],
@@
In the remainder of this section, we proceed by discussing the consequences a negative inductance on the dispersion relation. As already stated, for small $k$ we thus get an overtilted linear dispersion relation in both models, see Figs. 3a) and 3b). As the slope of the dispersion relation now has the same sign for both positive and negative $k$, we get signals propagating in the same direction. If we now join regions with regular ($u>|v|$) and overtilted ($u<|v|$) dispersion relation, we could indeed engineer analog wormholes. But before we embark on that, we need to discuss in detail the behaviour of the dispersion relations, Eq. (III), for large $k$. Namely, when considering the full Brillouin zone (i.e., for $-\pi<k<\pi$), we see that the two models (nearest versus next-nearest neighbour) differ significantly. In particular, for the nearest neighbor coupling, we observe that outside the range $-k_{0}<k<k_{0}$ the dispersion relation becomes complex, see Fig. 3a), where $k_{0}$ is determined by the condition
@@
To proceed, we note that in the 1D chain models we consider, there are ways to stabilize the system in spite of negative inductances, in line with the discussion in Sec. II. Indeed, the gyrators play the pivotal role for stabilization: in the Hamiltonians of Eqs. (III) and (III), when we expand the term in the first line $\sim E_{C}$, we get an effective inductive element due to a nonzero $G$, $\sim E_{C}G^{2}(\phi_{j+1}-\phi_{j-1})^{2}$. This is an effective inductive contribution due to the gyrator, which is guaranteed to be positive even when $E_{L}<0$ (as long as $E_{C}>0$). But as already foreshadowed in Sec. II.2, this effective inductive contribution couples next-nearest neighbour sites. Hence, a negative next-nearest inductor (Eq. (III)) can be exactly compensated by a positive next-nearest neighbour contribution of the gyrator. The negative nearest neighbour inductor model (Eq. (III)) on the other hand cannot exactly be compensated by the gyrators, hence the presence of the exceptional point.
@@
Let us now take a bit of more formal perspective on the above observation for generic wormhole simulations, which will allow us to demonstrate one of the central results of this work, see also item (iv) in Sec. II.2. Namely, there is a simple topological connection between overtilted spectra, Brillouin zone sizes, and exceptional points. For discrete lattice implementations the dispersion relation is always $2\pi$ periodic in $k$. Crucially, an overtilted spectrum means that the energies cross from positive to negative values at the transition points $k=0,2\pi,\ldots$. If the spectrum is real for all $k$, then periodicity in momentum space always guarantees that an overtilted spectrum near a certain value of $k$ must have an overtilted partner at another value of $k$ with opposite tilt direction, as a periodic spectrum must traverse the zero-energy line an even number of times. Consequently, for any lattice implementation, it seems unavoidable that what appears like a black hole horizon near one value of $k$, must necessarily have a while hole horizon partner at another value of $k$. This fact has already been pointed out in the context of analog black-hole simulations in tilted Weyl semimetal structures [25]. Our next-nearest neighbour model confirms this expectation, where the overtilt point at $k\approx 0$ has an inverted partner point at $k\approx\pm\pi$, see Fig. 3 b). These two $k$ points can be individually addressed when preparing, e.g., Gaussian wave packets centered either around $k\approx 0$ (wave packets that are smooth in $j$ space) and $k\approx\pm\pi$ (wave packets with alternating even-odd signs in $j$ space).
@@
Crucially, the nearest-neighbour model provides an intriguing loop hole: here, the overtilt at $k\approx 0$ has no partner point with inverted tilt, very simply because the dispersion relation satisfies the periodicity constraint in $k$ by taking, in a sense, a “detour” in the complex plane – by courtesy of the exceptional point. Consequently, the spectrum crosses the zero energy line only once. We are thus able to conclude that there is after all a case, where black and while hole horizons can be created independently in a lattice. This comes at the cost of a spontaneous instability within the overtilted region, which, importantly, is present even in the absence of an event horizon. This finding will be of great importance when interpreting the origin of radiation in the presence of event horizons in Sec. V below.

### v2

@@ Abstract
There has been considerable effort to simulate quantum gravity features in solid state systems, such as analog black holes or wormholes. However, superconducting circuits have so far received only limited attention in this regard. Moreover, for quantum superpositions of spacetime geometries – a highly contentious notion within the quantum gravity community itself – there currently exist no solid state blueprints. We here show how quantum circuit hardware can implement a variety of classical and quantum spacetime geometries on lattices, by both using established circuit elements and introducing new ones. We demonstrate the possibility of a metric sharply changing within a single lattice point, thus entering a regime where the spacetime curvature itself is trans-Planckian, and the Hawking temperature ill-defined. In fact, our approach suggests that stable, thermal event horizons are incompatible with strictly discrete lattice models. We thus propose to directly probe the evaporation of a wormhole by tracking the accumulation of charge and phase quantum fluctuations over short time scales, which are a robust signature even in the presence of a dissipative environment. Moreover, we present a loop-hole for the typical black/white hole ambiguity in lattice simulations: the existence of exceptional points in the dispersion relation allows for the creation of pure black (or white) hole horizons –- at the expense of radically changing the interior wormhole dynamics. Finally, based on multistable Josephson junctions, we introduce the notion of quantum inductors: circuit elements that can be prepared in a superposition of different inductance values. Such inductors realise regions with a quantum superposition of analog spacetime. The entanglement of signals with the quantum spacetime can be probed by a type of delayed-choice experiment.
@@ I Introduction
Unifying quantum mechanics and gravity poses several theoretical [1, 2, 3] and experimental challenges [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], which motivates studying analogs of relativistic quantum effects in experimentally accessible systems [15]. Several ideas have therefore been proposed for simulating gravitational effects in solid-state systems – for example, holographic ideas borrowed from string theory [16, 17, 18, 19, 20, 21], or cosmological particle creation [22, 23, 24, 25, 26, 27], or more directly, using Unruh’s proposal of "sonic black holes" [28] as an inspiration to simulate black holes in labs. Along these lines, apparent event horizons and the resulting Hawking radiation have been studied on various platforms  [29, 30, 31, 32, 33, 34, 35, 36, 37], most notably in ultracold atomic gases [38, 39, 40, 41, 42].
@@
There remain however a number of open questions. First, so far for all solid state implementations the metric can change only over a finite “healing length”. If the curvature at the horizon is too weak, the resulting small Hawking temperature may thus be below the threshold of the system’s intrinsic temperature [43] – an obstacle that (at least up to now) seems to have only been overcome for cold atom simulators [41]. Relatedly, systems with an apparent event horizon generically do not have a well-defined ground state, such that any coupling to the environment leads to an instability, making it extremely challenging to distinguish intrinsic radiation due to the horizon (which would actually simulate aspects of a universe with nontrivial spacetime metric) and radiation due to environment induced relaxation. In addition, even in the absence of environment, the closed system can in some platforms be spontaneously unstable [38, 39]. Furthermore, lattice realizations (which ultimately concerns all solid state systems) add a number of interesting but challenging facets, such as a natural (though possibly artifical) resolution of the trans-Planckian problem [44, 45, 46, 47], and most notably, the fact that any event horizon has both black and white hole character, since the dispersion relation in a periodic Brillouin zone must necessarily cross zero at least twice [31]. Finally, within the quantum gravity community, there also exists the highly contentious idea of superpositions of spacetime geometries, originating from attempts at quantizing the metric tensor [48, 49]. In particular, recent theoretical proposals suggest experiments that examine quantum gravitational effects caused by a superposition of different mass configurations [50, 51, 52, 53, 54, 55, 56, 57]. To the best of our knowledge, there currently exist no feasible blueprints for solid state systems capable of simulating quantum superpositions of spacetime geometries.
@@
Superconducting circuits are one of the prime candidates for building large-scale quantum hardware [58, 59, 60]. The behaviour of these circuits and their interactions with quantised electromagnetic fields in the microwave range are described by circuit quantum electrodynamics (cQED), a formalism that reduces circuits to lumped elements whose phase and charge are canonically conjugate [61, 62, 63, 64, 65, 66, 67, 68, 69]. The toolbox of superconducting circuits includes nonlinear elements such as Josephson junctions and linear elements such as capacitors, inductors, and gyrators – a nonreciprocal multi-port element [70, 71, 72, 73, 74, 75]. The same nonreciprocal behaviour also occurs due to topological transitions in the transport degrees of freedom of multiterminal junctions [76, 77, 78, 79, 80, 81]. Over the course of decades, it has been shown that networks of superconducting circuits can give rise to a large number of physical phenomena and imitate various quantum field theories from other domains [82, 83, 84, 85, 86, 87, 88, 89, 90, 34, 91]. However, the exploration of analog gravity phenomena has so far been limited to Hawking-like radiation in solitons of Josephson junction arrays [32, 33, 92, 93, 94, 95, 34, 35] which leave only little control over the shape of the spacetime geometry, or recreation of a limited class of curved spacetimes by means of flux control of SQUID arrays [36, 37].
@@
Moreover, we find an instructive loop-hole to the aforementioned ambiguity of black- versus white hole event horizons in lattice systems. Namely, the inductive coupling can go either via nearest or next-to-nearest neighbour nodes. For the former, the dispersion relation inside the wormhole region exhibits an exceptional point, such that the eigenspectrum no longer crosses zero energy twice, but instead takes a detour on the complex plane to satisfy periodicity. We thus realize lattice versions of event horizons which are definitely either black- or white hole, but not both. However, this feature radically changes the dynamics in the interior of the wormhole: instead of an evaporation from the horizons outwards the entire wormhole interior evaporates everywhere immediately after the quench.
@@
Finally, we go beyond the simulation of classical apparent horizons, and demonstrate that the cQED platform suggested here allows for the implementation of superpositions of spacetime geometries. This idea draws from the recent development of multistable Josephson junctions [100, 101, 102], i.e., junctions whose energy-phase relationship has more than one local minimum. Specifically, the engineering of multiple minimas with different local curvature creates a type of quantum inductor – an element that can be prepared in a superposition of different inductance values. When creating a chain of maximally entangled quantum inductors, we propose to create a superposition of different spacetime geometries. A classical signal traversing this region propagates in a superposition of different velocities with respect to the lab frame, and thus entangles with the quantum spacetime. The entanglement can be probed by a delayed-choice type of experiment, where a Hadamard gate is applied on the quantum inductor array after traversal of the quantum spacetime region, leading to a specific interference pattern.
@@
In an extended outlook, we show that in particular the idea of quantum superpositions of spacetime geometries harbors a great potential for follow-up studies, where one could endow the spacetime itself with a nonzero Hamiltonian, thus working towards much more complex simulations of analog quantum gravity, where both the field moving in curved analog spacetime and the spacetime itself have a quantum coherent and strongly correlated dynamics. Moreover, while we mostly consider long chains in the main part of this work, as they provide a clean interpretation of the effects in the context of curved spacetime, we are aware that they may be challenging to realize experimentally. In the outlook, we therefore also propose experiments on only few lattice points as proofs of principle, which already contain much of the pertinent phenomenology present in long chains.
@@
This paper is organized as follows, section II is a brief review of scalar field theories on curved spacetime, and a summary of our main accomplishments of this work. In section III we introduce and study the two circuits that will host the apparent horizons that we wish to investigate, we especially focus on the different effect that negative inductances have on the dispersion relations of these two circuits anticipating the differences between the horizons in them. Section IV is a brief detour, where we expand upon the idea of using flux quench on a Josephson junction to get a negative inductance, and in section V we study and characterize the horizons in the two aforementioned circuits, by looking at the time evolution of quantum fluctuations of the conjugate charge and the phase difference. In section VI we hypothesize on the long time fate of the analog horizons and argue that the effect of this intrinsic evaporation is easily distinguishable from the effect of coupling to the environment. Finally, in section VII we present our proposal for observing quantum superpositions of two spacetime geometries using multi-stable Josephson junctions.
@@ II.2 Simulation with circuit networks
As we see in Eq. (2), for a system to emulate a field moving on a nontrivial spacetime, we need interactions that provide the terms $\sim\dot{\phi}^{2}$$\sim\dot{\phi}\,\partial_{x}\phi$, and $\sim(\partial_{x}\phi)^{2}$. Quantum circuit theory [62, 64, 71, 69] readily provides elements that (in a certain continuum limit) provide such interactions. For a superconducting node $j$, a finite capacitance ($C_{j}$) provides to the Lagrangian $\mathcal{L}$ an energy contribution of the form
@@
    we show how to realize negative inductances via transient flux drive of Josephson junctions, and thus create overtilted dispersion relations and analog (apparent) event horizons.
This particular implementation of negative inductances via a flux quench is first of all of fundamental interest, as (contrary to other realizations [38, 42]) we do not require a continuous non-equilibrium drive to maintain the event horizons. As we detail further below, the absence of a well-defined ground state is merely the result of an approximation valid at short times. The exact quantum field theory retains a ground state at all times, which will ultimately be reached again via relaxation processes, thus potentially offering the possibility of a complete wormhole evaporation simulation emerging naturally from the hardware. Moreover, the setup we propose allows for a radical change of the spacetime geometry over only a few lattice sites (and as we show in a moment, even over a single lattice bond). With the energy scales of the local charging energy given by $E_{C}$ and when using Josephson junctions as inductive elements $E_{J}$, we would thus (at least in principle) be able to reach Hawking temperatures (as defined above) on the order of $\sim\sqrt{E_{C}E_{J}}$ which (taking as a typical transmon qubit frequency the value of $1\text{GHz}\sim 10\text{GHz}$) can yield up to $0.1\text{K}\sim 1\text{K}$ (which is above typical cryogenic temperatures).
@@
At any rate, some of the lattice-specific stability issues can be appreciated on the general level here, others will be shown subsequently for a specific model. On a general footing, we see in Eq. (2) that any (constant or spatially dependent) local tilt $v$ (which in our simulation would be implemented by constant or spatially varying gyrator parameters $G_{j}$) must equally appear in the $(\partial_{x}\phi)^{2}$ term, as the corresponding matrix element $g^{11}$ ($\mu,\nu=1$ is the space-space component) is proportional to $v^{2}-u^{2}$. In other words, a finite nonreciprocity in the circuit needs to be “countered” by a corresponding inductive interaction, in order to realize actions of the form of Eq. (2). Consequently, we would need an according spatially dependent tuning of the local inductance profile $L_{j}$ which matches exactly with the values of $G_{j}$. But first of all, such a perfect fine tuning is realistically not possible, such that the condition $g=-1$ is in general only approximately, but never exactly, fulfilled. Hence, the mapping onto Eq. (5) is likewise not exact, and we cannot guarantee a perfect causal decoupling across the horizon, and the closed system might become spontaneously unstable over already short time scales. In fact, below we refrain from a simulation with constant metric determinant $g$ (in fact, we will modify the inductances rather than the gyrator values), such that the systems we consider do not map to Eq. (5) in the continuum limit. Furthermore, with the Hawking temperature being defined (at least in continuous field theories) as the derivative of the metric (i.e., the curvature) at the event horizon, there emerges the curious question of what happens to the observables of the system when considering a very sharp (discontinuous) change of the metric at the horizon, and how such a diverging curvature is regularized. In what follows we will provide an answer to that question for lattice simulators.
@@
Thus, we see that in order to have the correct counter term in the $g^{11}$ component for the lattice, we actually do not need nearest neighbour inductive coupling [as indicated in Eq. (8)] but instead next-to-nearest neighbour coupling, $\sim(\phi_{j+1}-\phi_{j-1})^{2}$. As further detailed below, because of this fundamental feature of the gyrator element, nearest and next-to-nearest neighbour inductive couplings have fundamentally different stability properties. In particular,
@@
The breaking of charge conservation (which is physical as it corresponds to leakage to ground) effectively provides a mass term in the theory. In fact, we will include a (finite but very small) mass term in the following, but for a different reason: it allows us to deal with the otherwise problematic zero mode in the boson Hamiltonian.
@@ III Setup
We here show with two example circuit models the capability of circuit arrays to engineer nontrivial dispersion relations with analog wormholes. The first proposed circuit consists of an array of $J$ nodes (with periodic boundary conditions, i.e., $J+1=1$), interconnected via Josephson junctions in parallel with 3-port gyrators, as depicted in Fig. 1a). The complete Lagrangian of the circuit will consist of three parts, namely a capacitive Lagrangian in Eq. (7), a gyrator Lagrangian in Eq. (9) and a term for the Josephson junctions $\mathcal{L}_{J}=\sum_{j=1}^{J}E_{J}\cos\left(\phi_{j+1}-\phi_{j}+\phi_{\text{ext},j}\right)$. Here, $E_{J}$ is the Josephson energy and $\phi_{\text{ext},j}$ is the external flux coupling to each Josephson junction. This flux, and its time-dependent control, will play a pivotal role in what follows.
@@
As stated in the previous section, the stability of circuits where Josephson junctions connect nearest neighbours nodes and the ones where Josephson junctions couple next nearest neighbour nodes is very different. Therefore, to illustrate their differences the second circuit we study has next nearest neighbour connections via Josephson junctions, see Fig. 1b). The Hamiltonian for such a circuit is
@@
For small $k$, the system exhibits a linear dispersion relation, [see Figs. 2a) and 2b)],
@@
In the remainder of this section, we proceed by discussing the consequences a negative inductance has on the dispersion relation. As already stated, for small $k$ we thus get an overtilted linear dispersion relation in both models, see Figs. 3a) and 3b). As the slope of the dispersion relation now has the same sign for both positive and negative $k$, we get signals propagating in the same direction. If we now join regions with regular ($u>|v|$) and overtilted ($u<|v|$) dispersion relation, we could indeed engineer analog wormholes. But before we embark on that, we need to discuss in detail the behaviour of the dispersion relations, Eq. (III), for large $k$. Namely, when considering the full Brillouin zone (i.e., for $-\pi<k<\pi$), we see that the two models (nearest versus next-nearest neighbour) differ significantly. In particular, for the nearest neighbor coupling, we observe that outside the range $-k_{0}<k<k_{0}$ the dispersion relation becomes complex, see Fig. 3a), where $k_{0}$ is determined by the condition
@@
To proceed, we note that in the 1D chain models we consider, there are ways to stabilize the system in spite of negative inductances, in line with the discussion in Sec. II. Indeed, the gyrators play the pivotal role for stabilization: in the Hamiltonians of Eqs. (III) and (III), when we expand the term in the first line $\sim E_{C}$, we get an effective inductive element due to a nonzero $G$, $\sim E_{C}G^{2}(\phi_{j+1}-\phi_{j-1})^{2}$. This is an effective inductive contribution due to the gyrator, which is guaranteed to be positive even when $E_{L}<0$ (as long as $E_{C}>0$). But as already foreshadowed in Sec. II.2, this effective inductive contribution couples next-nearest neighbour sites. Hence, a negative next-nearest inductor [Eq. (III)] can be exactly compensated by a positive next-nearest neighbour contribution of the gyrator. The negative nearest neighbour inductor model [Eq. (III)] on the other hand cannot exactly be compensated by the gyrators, hence the presence of the exceptional point.
@@
Let us now take a bit of a more formal perspective on the above observation for generic wormhole simulations, which will allow us to demonstrate one of the central results of this work, see also item (iv) in Sec. II.2. Namely, there is a simple topological connection between overtilted spectra, Brillouin zone sizes, and exceptional points. For discrete lattice implementations the dispersion relation is always $2\pi$ periodic in $k$. Crucially, an overtilted spectrum means that the energies cross from positive to negative values at the transition points $k=0,2\pi,\ldots$. If the spectrum is real for all $k$, then periodicity in momentum space always guarantees that an overtilted spectrum near a certain value of $k$ must have an overtilted partner at another value of $k$ with opposite tilt direction, as a periodic spectrum must traverse the zero-energy line an even number of times. Consequently, for any lattice implementation, it seems unavoidable that what appears like a black hole horizon near one value of $k$, must necessarily have a white hole horizon partner at another value of $k$. This fact has already been pointed out in the context of analog black-hole simulations in tilted Weyl semimetal structures [31]. Our next-nearest neighbour model confirms this expectation, where the overtilt point at $k\approx 0$ has an inverted partner point at $k\approx\pm\pi$, see Fig. 3 b). These two $k$ points can be individually addressed when preparing, e.g., Gaussian wave packets centered either around $k\approx 0$ (wave packets that are smooth in $j$ space) and $k\approx\pm\pi$ (wave packets with alternating even-odd signs in $j$ space).
@@
Crucially, the nearest-neighbour model provides an intriguing loop hole: here, the overtilt at $k\approx 0$ has no partner point with inverted tilt, very simply because the dispersion relation satisfies the periodicity constraint in $k$ by taking, in a sense, a “detour” in the complex plane – by courtesy of the exceptional point. Consequently, the spectrum crosses the zero energy line only once. We are thus able to conclude that there is after all a case, where black and white hole horizons can be created independently in a lattice. This comes at the cost of a spontaneous instability within the overtilted region, which, importantly, is present even in the absence of an event horizon. This finding will be of great importance when interpreting the origin of radiation in the presence of event horizons in Sec. V below.